OpenStudy (anonymous):

simplify 14-7x over x-2

OpenStudy (anonymous):

Brad first factor out 7 from 14-7x

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

14 - x?

OpenStudy (anonymous):

-7(x-2) over x-2 -7

OpenStudy (anonymous):

when you factor out something, all you're doing is taking a constant that is a fraction of the variable and constants themselves for example (2x+2)=2(x+1) if i multiply this back again you get 2x+2 in your case 7(14-x)=98-7x which doesn't equal your original 14-7x

OpenStudy (anonymous):

@asifalibabar , unless you want to get hounded by a bunch of moderators, i suggest you don't simply give them the answer and allow the person you're helping to do it themselves.

OpenStudy (anonymous):

it's in the CoC. I've given more of an answer you have and was contacted by one so...

OpenStudy (anonymous):

But when u'll multiply this -7(x-2) u'll get this 14-7x

OpenStudy (anonymous):

\[\frac{14-7x}{x-2} \] \[\frac{7(2-x)}{-1(2-x)}\] \[ = -7 \]